OpenStudy (anonymous):

when balancing chemical equations , the number that is manipulated is the ? A. subscipts B.coeficient C.expontent D.atomic number

OpenStudy (anonymous):

@phi @hba @SWAG

OpenStudy (anonymous):

B) The coefficients indicate the molar relation between reactants and products For example \[ \text{Fe} + \text{Cl}_2 = \text{FeCl}_3\] is a the chemical equation for the reaction of iron(III) and chlorine into ferric chloride (iron(III) chloride ), but it is not balanced. The "number of chlorines" on the left does not equal the "number of chlorines" on the right. \[\begin{array} \text{Fe} + \text{Cl}_2 & = & \text{FeCl}_3 \\ \hline \\ \text{Fe}=1 & & \text{Fe}=1 \\ \text{Cl} = 2 & & \text{Cl}=3 \end{array}\] To balance the equation: \[N \ \text{Fe} + M \ \text{Cl}_2 = N \ \text{FeCl}_3\] \[ \text{Cl} \rightarrow \left| \begin{array}{l} 2M=3N \\ \ \\ \frac{M}{N}=\frac{3}{2} \\ \ \\ M=3 ; N=2 \end{array} \right. \] So \[\begin{array} \ 2 \text{Fe} + 3\text{Cl}_2 & = & 2\text{FeCl}_3 \\ \hline \\ \text{Fe}=2 & & \text{Fe}=2 \\ \text{Cl} = 6 & & \text{Cl}=6 \end{array}\] As opposed to: If you changed the subscripts, it would indicate that you were changing the chemical composition of the atoms and molecules involved in the reaction. If you changed the exponents, it would indicate that you were changing the charge of the atoms and molecules in the reaction. If you changed the atomic number, it would indicate that you were changing the internal nuclear structure of the atoms involved in the reaction.